### v2

# Gibbs sampling, exact sampling, and distribution function evaluation for the generalized inverse Gaussian distribution

Abstract
The generalized inverse Gaussian, denoted $\mathrm{GIG}(p,a,b)$, is a flexible family of distributions that includes the gamma, inverse gamma, and inverse Gaussian distributions as special cases. In addition to its applications in statistical modeling and its theoretical interest, the GIG often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. This article introduces two mixture representations for the GIG: one that expresses the distribution as a continuous mixture of inverse Gaussians and another that reveals a recursive relationship between GIGs with different values of $p$. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This simple Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to algorithms for exact, rejection-free sampling as well as CDF evaluation for the GIG with half-integer $p.$ We highlight computational examples from the literature where these new algorithms could be applied.
1 Introduction
The generalized inverse Gaussian, which we denote $\mathrm{GIG}(p,a,b)$, is a three-parameter, absolutely continuous distribution supported on $\mathbb{R}_{+}$. It is a rich family of distributions that encompasses the gamma distribution (if $b=0$), the inverse gamma distribution (if $a=0$), and the inverse Gaussian distribution (if $p=-1/2$). The parameters of the $\mathrm{GIG}(p,a,b)$ are $p\in\mathbb{R}$, $a>0$, and $b>0$, and its probability density function (PDF) is
[…]
where $K_{p}(\cdot)$ is the modified Bessel function of the second kind [1]. The moments of the $\mathrm{GIG}(p,a,b)$ involve Bessel functions but are available in closed form (see, for example, [22]).
The generalized inverse Gaussian was first proposed by Halphen in the 1940s for analyzing hydrological data [15, 30]. It regained popularity in the mid to late 1970s, when a series of articles studied its properties in detail (see, for example, [5], [8], and [14]). For an account of known properties and characterizations of the $\mathrm{GIG}(p,a,b)$, we refer the reader to the monograph [22] and the review article [24].
The GIG$(p,a,b)$ often arises in computational statistics, especially in Markov chain Monte Carlo (MCMC) algorithms for posterior inference. For example, it emerges as a full conditional distribution for scale parameters in normal models. Sections 2 and 3 describe or reference a variety of applications where it is necessary to simulate random draws from the $\mathrm{GIG}(p,a,b)$ or evaluate its cumulative distribution function (CDF).
The generalized inverse Gaussian is challenging to simulate from or evaluate the CDF of. Several articles have been written on generating random draws from the $\mathrm{GIG}(p,a,b)$, including [11], [25], [19], [12], and [32]. Recently, [31] derived expressions for the CDF of the generalized inverse Gaussian for half-integer $p$.
This article introduces two novel mixture representations of the generalized inverse Gaussian. One expresses the $\mathrm{GIG}(p,a,b)$ as a continuous mixture of inverse Gaussians, while the other reveals a relationship between GIGs with different values of $p$. The former representation forms the basis for a data augmentation scheme that leads to a geometrically ergodic Gibbs sampler for the GIG. This simple Gibbs sampler, which alternates between gamma and inverse Gaussian conditional distributions, can be incorporated within an encompassing MCMC algorithm when simulation from a GIG is required. The latter representation leads to the first algorithm in the literature for exact, rejection-free sampling from the GIG with half-integer $p$ as well as an algorithm for CDF evaluation in that same setting. The algorithms are included in the article and can be easily implemented in standard programming languages. All proofs appear in the Supplementary Material.
1.1 Notation
We use the notation $\text{InvGauss}(\mu,\lambda)$ for the inverse Gaussian distribution with mean parameter $\mu$ and shape parameter $\lambda$, $\text{Gamma}(\alpha,\beta)$ for the gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$, $\text{Exp}(\theta)$ for the exponential distribution with rate parameter $\theta$, $\text{Normal}(\mu,\sigma^{2})$ for the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and $\text{Unif}(0,1)$ for the uniform distribution between zero and one. We denote the PDF of the $\mathrm{GIG}(p,a,b)$ evaluated at $x$ by $\mathrm{GIG}(x\mid p,a,b)$. We use analogous notation for the PDFs of other distributions.
2 Data-augmented Gibbs sampler
In this section, we introduce a representation of the GIG$(p,a,b)$ as a continuous mixture of inverse Gaussian distributions. The representation is useful for deriving a data-augmented Gibbs sampler for which the stationary distribution of one of the marginals is GIG$(p,a,b)$ (Algorithm 1). We then reference examples from the literature where the GIG appears as a full-conditional distribution in Gibbs sampling algorithms. Example 1 shows how Algorithm 1 can be nested within an encompassing MCMC algorithm when simulation from a GIG is required. Finally, we provide theoretical support for the Gibbs sampler described in Algorithm 1. In particular, Proposition 3 establishes that it is geometrically ergodic.
Proposition 1.
The PDF of the $\mathrm{GIG}(p,a,b)$ for $p\neq-1/2$ can be written as a continuous mixture of inverse Gaussian distributions:
$f_{X}(x)=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y,$
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF. The parameters of the distributions depend on whether $p<-1/2$ or $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)$
$f_{Y}(y)$ $=\frac{1}{\mathcal{Z}_{0}}\,y^{-(p+3/2)}\,\exp\left\{-\sqrt{b(a+2y)}\right\}$
$\mathcal{Z}_{0}$ $=\frac{b^{(p+1)/2}\,\sqrt{2}\,\Gamma(-p-1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}}.$
If $p>-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)$
$f_{Y}(y)$ $=\frac{1}{\mathcal{Z}_{1}}\,\frac{y^{p-1/2}}{\sqrt{b+2y}}\,\exp\left\{-\sqrt{a(b+2y)}\right\}$
$\mathcal{Z}_{1}$ $=\frac{b^{p/2}\,\sqrt{2}\,\Gamma(p+1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}}.$
By Proposition 1, we can define a joint PDF $f_{X,Y}(x,y)=f_{X\mid Y}(x)f_{Y}(y)$ whose marginal $f_{X}(x)$ is $\mathrm{GIG}(x\mid p,a,b)$. We have established that $X\mid Y=y$ is inverse Gaussian. The Proposition below states that $Y\mid X=x$ is gamma-distributed.
Proposition 2.
Let $f_{X\mid Y}(x)$ and $f_{Y}(y)$ be as defined in Proposition 1. If $p<-1/2$, then $Y\mid X=x\sim\mathrm{Gamma}(-(p+1/2),x)$. If $p>-1/2$, then $Y\mid X=x\sim\mathrm{Gamma}(p+1/2,1/x)$.
We can construct a data-augmented Gibbs sampler by sampling iteratively from the conditional distributions $X\mid Y=y$ and $Y\mid X=x$ (see Algorithm 1). Only the draws from $X\mid Y=y$ are relevant if we are interested in sampling from the GIG$(p,a,b)$. The case $p=-1/2$ is not included in Propositions 1 and 2, but the $\mathrm{GIG}(-1/2,a,b)$ is simply $\mathrm{InvGauss}(\sqrt{b/a},b)$ [22].
There are efficient algorithms to generate random draws from the inverse Gaussian and gamma distributions. To generate inverse Gaussian draws, one can use the algorithm described in [28], whereas for the gamma distribution, one can use standard algorithms such as those proposed in [2].
Algorithm 1 Data-augmented composition Gibbs sampler for $\mathrm{GIG}(p,a,b)$
$p\in\mathbb{R},a>0,b>0,n_{\mathrm{sim}}\in\{1,2,\,...\}$
$x\leftarrow\text{vector}(\text{length}=n_{\mathrm{sim}})$
if $p<-1/2$ then
     $y\leftarrow-(p+1/2)\sqrt{b/a}$
     for $i$ in 1 to $n_{\mathrm{sim}}$ do
         $x[i]\sim\mathrm{InvGauss}(\sqrt{b/(a+2y)},b)$
         $y\sim\mathrm{Gamma}(-(p+1/2),x)$
     end for
else if $p=-1/2$ then
     for $i$ in 1 to $n_{\mathrm{sim}}$ do
         $x[i]\sim\mathrm{InvGauss}(\sqrt{b/a},a)$
     end for
else
     $y\leftarrow(p+1/2)\sqrt{b/a}$
     for $i$ in 1 to $n_{\mathrm{sim}}$ do
         $x[i]\sim\mathrm{InvGauss}(\sqrt{(b+2y)/a},b+2y)$
         $y\sim\mathrm{Gamma}(p+1/2,1/x)$
     end for
end if
return $x$
The GIG$(p,a,b)$ often arises as a full conditional distribution in Gibbs sampling algorithms. There are countless examples from the literature on global-local shrinkage priors [3, 33, 7, 16], graphical models [23, 29], stochastic volatility models [6], Bayesian nonparametrics [13], smoothing splines [27], and other areas. The Gibbs sampler described in Algorithm 1 can be embedded within an encompassing MCMC algorithm when simulation from a GIG$(p,a,b)$ is required. The following example offers a simple stylized illustration.
Example 1.
Suppose that we observe independent and identically distributed data $y_{i}\mid\mu,\sigma^{2}\sim\mathrm{Normal}\left(\mu,\sigma^{2}\right)$, $i\in\{1,2,\,...\,,n\}$, with $\mu$ and $\sigma^{2}$ unknown. We assign independent priors to $\mu$ and $\sigma^{2}$ with $\mu\sim\mathrm{Normal}\left(\theta_{0},\tau_{0}^{2}\right)$ and $\sigma^{2}\sim\mathrm{GIG}\left(p_{0},a_{0},b_{0}\right).$ Letting $\mathbf{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$ and $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i},$ we can simulate from the posterior distribution of $\left(\mu,\sigma^{2}\right)$ by iterating between the full conditional distributions
$\mu\mid\sigma^{2},\mathbf{y}$ $\sim\mathrm{Normal}\left(\theta_{n},\tau_{n}^{2}\right)$
$\sigma^{2}\mid\mu,\mathbf{y}$ $\sim\mathrm{GIG}\left(p_{n},a_{n},b_{n}\right),$
where
$\theta_{n}=\left(\frac{n}{\sigma^{2}}+\frac{1}{\tau_{0}^{2}}\right)^{-1}\left(\frac{n}{\sigma^{2}}\bar{y}+\frac{1}{\tau_{0}^{2}}\theta_{0}\right),\quad\tau_{n}^{2}=\left(\frac{n}{\sigma^{2}}+\frac{1}{\tau_{0}^{2}}\right)^{-1},$
and
$p_{n}=p_{0}-\frac{n}{2},\quad a_{n}=a_{0},\quad b_{n}=b_{0}+\sum_{i=1}^{n}(y_{i}-\mu)^{2}.$
We can leverage the results of this section to derive a Gibbs sampler that does not require simulating directly from the GIG$(p,a,b)$. If $p_{n}<-1/2,$ we can introduce an auxiliary variable $\omega$ and iterate through the full conditional distributions
$\mu\mid\sigma^{2},\mathbf{y}$ $\sim\mathrm{Normal}\left(\theta_{n},\tau_{n}^{2}\right)$
$\sigma^{2}\mid\mu,\omega,\mathbf{y}$ $\sim\mathrm{InvGauss}\left(\sqrt{b_{n}/(a_{n}+2\omega)},b_{n}\right)$
$\omega\mid\mu,\sigma^{2},\mathbf{y}$ $\sim\mathrm{Gamma}\left(-(p_{n}+1/2),\sigma^{2}\right).$
If $p_{n}>-1/2,$ the full conditional distributions are
$\mu\mid\sigma^{2},\mathbf{y}$ $\sim\mathrm{Normal}\left(\theta_{n},\tau_{n}^{2}\right)$
$\sigma^{2}\mid\mu,\omega,\mathbf{y}$ $\sim\mathrm{InvGauss}\left(\sqrt{(b_{n}+2\omega)/a_{n}},b_{n}+2\omega\right)$
$\omega\mid\mu,\sigma^{2},\mathbf{y}$ $\sim\mathrm{Gamma}\left(p_{n}+1/2,1/\sigma^{2}\right).$
The Gibbs sampler in Algorithm 1 is geometrically ergodic, which is a property that quantifies the rate of convergence of the Gibbs sampler to its stationary distribution. It also ensures that inferences made with draws from the chain are well-behaved. In particular, if $\{X_{i}\}_{i=1}^{n}$ are draws from the chain and we want to estimate a function $f(X)$ with finite $\mathbb{E}[f(X)^{2+\varepsilon}]$ for some $\varepsilon>0$, geometric ergodicity ensures that $\sum_{i=1}^{n}f(X_{i})/n$ is approximately normal [10].
More formally, geometric ergodicity can be defined as follows (our exposition mirrors that in [21]). Let the $k$-step ahead transition kernel of the Markov chain be
$P^{k}[(x,y),A]=P[(X_{i+k},Y_{i+k})\in A\mid(X_{i},Y_{i})=(x,y)],$
where $A$ is a $P^{k}$-measurable set. In our context, $A\in\mathcal{B}(\mathbb{R}^{2}_{+})$, where (+$\mathcal{B}(\mathbb{R}^{2}_{+})$ are the Borel sets on +$\mathbb{R}^{2}_{+}$. Let $\pi(\cdot)$ be the probability measure of the stationary distribution of the chain. Then, geometric ergodicity requires the existence of a function $M(\cdot,\cdot)$ and a constant $t\in(0,1)$ so that
$\sup_{A\in\mathcal{B}(\mathbb{R}^{2}_{+})}|P^{k}[(x,y),A]-\pi(A)|\leq t^{k}M(x,y)$
for all $(x,y)$ in the support of $(X,Y)$, which in our case is +$\mathbb{R}^{2}_{+}$. In other words, the total variation distance between the $k$-step ahead kernel and the stationary distribution decreases exponentially fast in $k$ for all $(x,y)\in\mathbb{R}^{2}_{+}$.
We use Theorem 3.5 in [21] to establish that the Gibbs sampler described in Algorithm 1 is geometrically ergodic. Since the densities involved are all continuous, the proof amounts to showing that there exist functions $f,g:[0,\infty)\rightarrow[1,\infty)$ and constants $j,k,m,n>0$ with $jm<1$ so that
$\mathbbm{E}[f(X)\mid Y=y]\leq jg(y)+k,\qquad\mathbbm{E}[g(Y)\mid X=x]\leq mf(x)+n,$
with $C_{d}=\{y:g(y)\leq d\}$ compact for all $d>0$.
Proposition 3.
The data-augmented composition Gibbs sampler in Algorithm 1 is geometrically ergodic.
Algorithm 1 describes a composition Gibbs sampler because $X$ and $Y$ are always sampled in the same order. Random sequence scan and random scan Gibbs samplers can be defined in an analogous manner. In random sequence scan and random scan Gibbs samplers, the order in which $X$ and $Y$ are sampled is random and can change from iteration to iteration (see, for example, [21] for further details). The results in [21] imply that the analogous random sequence scan and random scan Gibbs samplers are also geometrically ergodic.
3 Exact sampling and CDF evaluation for half-integer $p$
The main results of this section reveal a relationship between GIGs with different values of $p.$ Proposition 4 leads to the first algorithm in the literature for exact, rejection-free simulation from the GIG$(p,a,b)$ for half-integer $p.$ Proposition 5 leads to a simple algorithm for evaluating the CDF of the GIG$(p,a,b)$ for half-integer $p.$ The results can be established using Theorem 2.4 of [20].
Proposition 4.
If $X\sim\mathrm{GIG}(p,a,b)$ and $p>1$, then $X=_{d}Y+E$ for independent random variables $Y$ and $E$ with
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2).$
The exact sampling algorithm for half-integer $p$ (Algorithm 2) is a direct consequence of Proposition 4. As described, the algorithm generates a single random number. To draw more than one number, one can wrap Algorithm 2 within a for loop.
Algorithm 2 rgig.half: Sampling from $\mathrm{GIG}(p,a,b)$ for half-integer $p$
$p\in\{\,...\,,-3/2,-1/2,1/2,3/2,\,...\},a>0,b>0$
if $p<0$ then
     $x\leftarrow 1/\mathrm{rgig.half}(-p,b,a)$
else if $p=1/2$ then
     $y\sim\mathrm{InvGauss}(\sqrt{a/b},a)$
     $x\leftarrow 1/y$
else
     $w\leftarrow K_{p-2}(\sqrt{ab})/K_{p}(\sqrt{ab})$
     $U\sim\mathrm{Unif}(0,1)$
     $E\sim\mathrm{Exp}(a/2)$
     if $U<w$ then
         $x\leftarrow\mathrm{rgig.half}(p-2,a,b)+E$
     else
         $x\leftarrow\mathrm{rgig.half}(p-1,a,b)+E$
     end if
end if
return $x$
The algorithm for evaluating the CDF of the GIG for half-integer $p$ (Algorithm 3) relies on the following recurrence formula, which is a consequence of Proposition 4.
Proposition 5.
Let $G_{p}(x)$ be the CDF of the $\mathrm{GIG}(p,a,b)$. For half-integer $p$ such that $|p|>1/2$, the following recurrence holds:
$G_{p}(x)=wG_{p-2}(x)+(1-w)G_{p-1}(x)-\exp\left\{\frac{-ax}{2}\right\}I_{p}(x),$
where
$I_{p}(x)=\frac{w(ab)^{(p-2)/2}}{2^{p-1}K_{p-2}(\sqrt{ab})}\Gamma\left(2-p,\frac{b}{2x}\right)+\frac{(1-w)(ab)^{(p-1)/2}}{2^{p}K_{p-1}(\sqrt{ab})}\Gamma\left(1-p,\frac{b}{2x}\right)$
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1].
The CDFs $G_{-1/2}$ and $G_{1/2}$ are both known. The former is the CDF of $\mathrm{InvGauss}(\sqrt{b/a},b)$. The latter can be evaluated with the CDF of an inverse Gaussian because if $X\sim\mathrm{GIG}(-p,b,a)$, then $1/X\sim\mathrm{GIG}(p,a,b)$ [22].
Proposition 5 suggests the implementation of a recursive algorithm that repeatedly calls the function to obtain $G_{p-2}$ and $G_{p-1}$ until reaching the base case. However, the bottom-up approach described in Algorithm 3 is substantially faster for moderate to large $p$ because it avoids redundant computations. The algorithm could be optimized further by storing the evaluations of the Bessel functions as they are computed. We do not pursue this in Algorithm 3 to simplify our exposition.
Algorithm 3 pgig.half: CDF of $\mathrm{GIG}(p,a,b)$ for half integer $p$
$x>0,p\in\{\,...\,,-3/2,-1/2,1/2,3/2,\,...\},a>0,b>0$
if $p<0$ then
     $1-\mathrm{pgig.half}(1/x,-p,b,a)$
else
     $v\leftarrow\mathrm{vector}(\text{length}=p+1/2)$
     $v[1]\leftarrow P[\mathrm{InvGauss}(\sqrt{a/b},b)>1/x]$
     if $\text{length}(v)>1$ then
         $w\leftarrow K_{-1/2}(\sqrt{ab})/K_{3/2}(\sqrt{ab})$
         $I\leftarrow\frac{w(ab)^{-1/4}\Gamma\left(1/2,\frac{b}{2x}\right)}{2^{1/2}K_{-1/2}(\sqrt{ab})}+\frac{(1-w)(ab)^{1/4}\Gamma\left(-1/2,\frac{b}{2x}\right)}{2^{3/2}K_{1/2}(\sqrt{ab})}$
         $k\leftarrow\exp\{-ax/2\}$
         $v[2]\leftarrow w\,\text{pgig.half}(x,-1/2,a,b)+(1-w)\,v[1]-kI$
         if $\text{length}(v)>2$ then
              for $i\text{ in }3:\text{length}(v)$ do
                  $w\leftarrow K_{i-5/2}(\sqrt{ab})/K_{i-1/2}(\sqrt{ab})$
                  $I\leftarrow\frac{w(ab)^{(i-5/2)/2}\Gamma\left(5/2-i,\frac{b}{2x}\right)}{2^{i-3/2}K_{i-5/2}(\sqrt{ab})}+\frac{(1-w)(ab)^{(i-3/2)/2}\Gamma\left(3/2-i,\frac{b}{2x}\right)}{2^{i-1/2}K_{i-3/2}(\sqrt{ab})}$
                  $v[i]\leftarrow w\,v[i-2]+(1-w)\,v[i-1]-kI$
              end for
         end if
     end if
     return $v[\text{length}(v)]$
end if
The base cases of the algorithms involve the inverse Gaussian distribution. Generating random draws from the inverse Gaussian and evaluating its CDF is straightforward. Both algorithms require the evaluation of the modified Bessel function of the second kind for half-integer orders of $p$. This can be done using well-known properties (see, for instance, [1]): (1) $K_{-\nu}(z)=K_{\nu}(z)$, (2) $K_{1/2}(z)=\sqrt{\pi/2z}\,\exp\{-z\}$, and (3) the recurrence formula $K_{\nu+1}(z)=2\nu K_{\nu}(z)/z+K_{\nu-1}(z).$
It is not uncommon that statistical methods require sampling from or evaluating the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$ [13] introduce a stick-breaking representation for the normalized inverse Gaussian process [26] whose practical value depends upon being able to efficiently simulate from the GIG for half-integer $p.$ In both [7] and [27], GIG$(p,a,b)$ full conditional distributions have half-integer $p$ parameters for the priors considered. [18] present a data augmentation scheme for models that include gamma functions that requires simulating from a sum of independent GIGs for which $p=-3/2.$ [31] describe several financial and actuarial applications in which one needs to evaluate the CDF of the $\text{GIG}(p,a,b)$ for half-integer $p.$
4 Conclusions
This article introduced representations of the generalized inverse Gaussian distribution that lead to simple algorithms for generating random draws and evaluating its CDF. As future work, it would be interesting to study if there are analogous representations for the matrix-variate version of the generalized inverse Gaussian [4, 9, 17], a probability distribution over symmetric positive-definite matrices that includes the Wishart and inverse Wishart as special cases.
Appendix A Proofs of Propositions
Proposition 1.
The PDF of the $\mathrm{GIG}(p,a,b)$ for $p\neq-1/2$ can be written as a continuous mixture of inverse Gaussian distributions:
$f_{X}(x)=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y,$
where $f_{X}(x)$ is the PDF of the $\mathrm{GIG}(p,a,b)$, $f_{Y}(y)$ is a PDF with support on $\mathbb{R}_{+}$, and $f_{X\mid Y}(x\mid y)$ is an inverse Gaussian PDF. The parameters of the distributions depend on whether $p<-1/2$ or $p>-1/2$. If $p<-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)$
$f_{Y}(y)$ $=\frac{1}{\mathcal{Z}_{0}}\,y^{-(p+3/2)}\,\exp\left\{-\sqrt{b(a+2y)}\right\}$
$\mathcal{Z}_{0}$ $=\frac{b^{(p+1)/2}\,\sqrt{2}\,\Gamma(-p-1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}}.$
If $p>-1/2$, then
$f_{X\mid Y}(x\mid y)$ $=\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)$
$f_{Y}(y)$ $=\frac{1}{\mathcal{Z}_{1}}\,\frac{y^{p-1/2}}{\sqrt{b+2y}}\,\exp\left\{-\sqrt{a(b+2y)}\right\}$
$\mathcal{Z}_{1}$ $=\frac{b^{p/2}\,\sqrt{2}\,\Gamma(p+1/2)\,K_{p}(\sqrt{ab})}{a^{p/2}\,\sqrt{\pi}}.$
Proof.
First, we find $f_{X\mid Y}(x\mid y)$. Along the way, we derive $f_{Y\mid X}(y\mid x)$, which will be useful for deriving a Gibbs sampler. We consider two cases: $p<-1/2$ and $p>-1/2$.
Case $p<-1/2$: Up to proportionality constants, the PDF $f_{X}(x)$ of the $\mathrm{GIG}(p,a,b)$ is
$f_{X}(x)\propto x^{p+1/2}\,\mathrm{InvGauss}(x\mid\sqrt{b/a},b).$
The term $x^{p+1/2}$ can be represented as follows
$x^{p+1/2}\underset{(p<-1/2)}{=}\frac{\int_{0}^{\infty}y^{-(p+3/2)}\,\exp\left(-xy\right)\,\mathrm{d}y}{\Gamma(-(p+1/2))}\,,$
so we can write
$f_{X}(x)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,\int_{0}^{\infty}y^{-(p+3/2)}\,\exp\left(-xy\right)\,\mathrm{d}y.$
If we define
$g(x,y)=\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right)\mathbbm{1}(x>0,y>0),$
then $g(x,y)$ is non-negative and continuous. Therefore, by Tonelli’s theorem, the double integral on +$\mathbbm{R}^{2}_{+}$ of $g(x,y)$ is equal to the iterated integral, which is finite. This implies that $g(x,y)$ can be normalized to be a joint PDF $f_{X,Y}(x,y)\propto g(x,y)$ whose marginal in $X$ is $f_{X}(x)$. By inspecting $g(x,y)$, we can easily deduce that
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{b/(a+2y)},b)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
Case $p>-1/2$: We use the same argument for $p<-1/2$ but with a different integral representation for $x^{p+1/2}$. In this case, we use the representation
$x^{p+1/2}=\frac{\int_{0}^{\infty}y^{p-1/2}\,\exp(-y/x)\,\mathrm{d}y}{\Gamma(p+1/2)}.$
We can rewrite the PDF of the $\mathrm{GIG}(p,a,b)$ as
$f_{X}(x)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,\int_{0}^{\infty}y^{p-1/2}\,\exp\left(-y/x\right)\,\mathrm{d}y.$
By Tonelli’s theorem, the joint PDF of $(X,Y)$ is proportional to
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y).$
The conditional distributions are
$X\mid Y=y$ $\sim\mathrm{InvGauss}(\sqrt{(b+2y)/a},b+2y)$
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Lastly, we find $f_{Y}(y)$. Again, we treat the cases $p<-1/2$ and $p>-1/2$ separately. If $p<-1/2$, then:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{\mathrm{InvGauss}(x\mid\sqrt{b/(a+2y)},b)}{\mathrm{Gamma}(y\mid-(p+1/2),x)}\,\mathrm{d}x$
$=\mathcal{Z}_{0}\,y^{p+3/2}\,\exp(b^{1/2}(a+2y)^{1/2}),$
where
$\mathcal{Z}_{0}=\frac{b^{(p+1)/2}2^{1/2}\Gamma(-p-1/2)K_{p}(a^{1/2}b^{1/2})}{a^{p/2}\pi^{1/2}}.$
If $p>-1/2$:
$f_{Y}(y)^{-1}$ $=\int_{0}^{\infty}\frac{f_{X}(x\mid y)}{f_{Y}(y\mid x)}\,\mathrm{d}x$
$=\int_{0}^{\infty}\frac{\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y)}{\mathrm{Gamma}(t\mid p+1/2,1/x)}\,\mathrm{d}x$
$=\mathcal{Z}_{1}\,y^{-(p-1/2)}\,(b+2y)^{1/2}\,\exp\left\{a^{1/2}(b+2y)^{1/2}\right\},$
where
$\mathcal{Z}_{1}=\frac{2^{1/2}\Gamma(p+1/2)K_{p}\left(a^{1/2}b^{1/2}\right)}{\pi^{1/2}(a/b)^{p/2}}.$
Proposition 2.
Let $f_{X\mid Y}(x)$ and $f_{Y}(y)$ be as defined in Proposition 1. If $p<-1/2$, then $Y\mid X=x\sim\mathrm{Gamma}(-(p+1/2),x)$. If $p>-1/2$, then $Y\mid X=x\sim\mathrm{Gamma}(p+1/2,1/x)$.
Proof.
We found these distributions in Proposition 1. They can be derived after inspecting the joint distributions (up to proportionality constants). If $p<-1/2$, then
$f_{X,Y}(x,y)\propto\mathrm{InvGauss}(x\mid\sqrt{b/a},b)\,y^{-(p+3/2)}\,\exp\left(-xy\right),$
from which it is straightforward to deduce that
$Y\mid X=x$ $\sim\mathrm{Gamma}(-(p+1/2),x).$
If $p>-1/2$, then
$f_{X,Y}(x,y)\propto y^{p-1/2}\exp\left\{-\sqrt{a(b+2y)}\right\}\,\mathrm{InvGauss}(x\mid\sqrt{(b+2y)/a},b+2y),$
which implies that
$Y\mid X=x$ $\sim\mathrm{Gamma}(p+1/2,1/x).$
Proposition 3.
The data-augmented composition Gibbs sampler in Algorithm 1 is geometrically ergodic.
Proof.
As we stated in the main text, we prove this result by checking that the conditions of Theorem 3.5 in [21] are satisfied. The proof amounts to showing that there exist functions $f,g:[0,\infty)\rightarrow[1,\infty)$ and constants $j,k,m,n>0$ with $jm<1$ so that
$\mathbbm{E}[f(X)\mid Y=y]\leq jg(y)+k,\qquad\mathbbm{E}[g(Y)\mid X=x]\leq mf(x)+n,$
with $C_{d}=\{y:g(y)\leq d\}$ compact for all $d>0$. We prove the results for $p<-1/2$ and $p>1/2$ separately.
Case $p<-1/2$: Define
$\theta$ $=3/(2b)$
$\mu$ $=3(1/2-p)/(2b)$
$f(x)$ $=1+\alpha(1/x-\theta)^{2},\qquad 0<\alpha<b^{2}/(p^{2}-1/4)$
$g(y)$ $=1+\beta(y-\mu)^{2},\qquad\alpha^{2}/b^{2}<\beta<\alpha/(p^{2}-1/4).$
The set $\{y:g(y)\leq d\}$ is compact for all $d>0$ and $f,g:[0,\infty)\rightarrow[1,\infty)$.
We find the conditional expectations of $f$ and $g$ and bound them. We start with the conditional expectation of $f$:
$\mathbbm{E}[f(X)\mid Y=y]$ $=1+\alpha\left\{\mathrm{Var}(1/X\mid Y=y)+\left[\mathbbm{E}(1/X\mid Y=y)-\theta\right]^{2}\right\}.$
Using well-known properties of the inverse Gaussian distribution, we find
$\mathrm{Var}(1/X\mid Y=y)$ $={(a+2y)^{1/2}}{b^{-3/2}}+{2}{b^{-2}}$
$\mathbbm{E}(1/X\mid Y=y)-\theta$ $=(a+2y)^{1/2}{b}^{-1/2}+b^{-1}-(3/2)b^{-1}.$
Rearranging terms, we obtain
$\mathbbm{E}[f(X)\mid Y=y]$ $=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha y/b$
$=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+2\alpha(y-\mu)/b$
$\leq 1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b+g(y),$
where the last inequality holds because $\beta\geq\alpha^{2}/b^{2}$. The inequality can be rewritten as
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=1$
$k$ $=1+9\alpha/(4b^{2})+\alpha a/b+2\alpha\mu/b.$
Both $j$ and $k$ are positive, which is required by Theorem 3.5 in [21]. Now, we find the conditional expectation of $g$ and bound it:
$\mathbbm{E}[g(Y)\mid X=x]$ $=1+\beta\left\{\mathrm{Var}(Y\mid X=x)+\left[\mathbbm{E}(Y\mid X=x)-\mu\right]^{2}\right\}$
$=1+\beta\left\{-(p+1/2)x^{-2}+\left[-(p+1/2)x^{-1}-\mu\right]^{2}\right\}$
$=1+\beta(1/2-p)/(4b^{2})+{\beta(p^{2}-1/4)}\left[1/x-\theta\right]^{2}$
$\leq 1+9\beta(1/2-p)/(4b^{2})+\beta(p^{2}-1/4)f(x)/\alpha.$
We can rewrite the inequality as
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=\beta(p^{2}-1/4)/\alpha$
$n$ $=1+9\beta(1/2-p)/(4b^{2}),$
where $m,n>0$, as required by Theorem 3.5 in [21]. It remains to show that $jm<1$. Given our choices of $\alpha$ and $\beta$,
$10<{\beta(p^{2}-1/4)}/{\alpha}<$ 1
and the interval $(\alpha^{2}/b,\alpha/(p^{2}-1/4))$ is nonempty. This completes the proof for this case.
Case $p>-1/2:$ Let $10<\gamma<$ 1 and
$\theta$ $=1/(2a)$
$\mu$ $=(p+1/2)/(2a)$
$f(x)$ $=1+\alpha(x-\theta)^{2},\qquad\alpha>0$
$g(y)$ $=1+\beta(y-\mu)^{2},\qquad\beta>\max[\alpha^{2}/(\gamma^{2}a),{\alpha}/{(p+1/2)^{2}}].$
Again, $\{y:g(y)\leq d\}$ is compact for all $d>0$ and $f,g:[0,\infty)\rightarrow[1,\infty)$. We bound the conditional expectations. We start with $f$:
$\mathbbm{E}[f(X)\mid Y=y]$ $=1+\alpha\left\{\mathrm{Var}(X\mid Y=y)+\left[\mathbbm{E}(X\mid Y=y)-\theta\right]^{2}\right\}$
$=1+\alpha\left\{{(b+2y)^{1/2}}{a^{-3/2}}+\left[(b+2y)^{1/2}a^{-1/2}-\theta\right]^{2}\right\}$
$=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+{2\alpha}(y-\mu)/a$
$\leq 1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a+\gamma g(y).$
The inequality holds because $\beta>\alpha^{2}/(\gamma^{2}a^{2})$. Therefore,
$\mathbbm{E}[f(X)\mid Y=y]$ $\leq jg(y)+k$
$j$ $=\gamma$
$k$ $=1+\alpha/(4a^{2})+\alpha b/a+{2\alpha}\mu/a,$
where $j$ and $k$ are positive, as required by Theorem 3.5 in [21]. Finally, we bound the conditional expectation of $g$:
$\mathbbm{E}[g(Y)\mid X=x]$ $=1+\beta\left\{\mathrm{Var}(Y\mid X=x)+[\mathbbm{E}(Y\mid X=x)-\mu]^{2}\right\}$
$=1+\beta\left\{(p+1/2)x^{2}+[(p+1/2)x-\mu]^{2}\right\}$
$=1+\beta(p+1/2)x^{2}+\beta(p+1/2)^{2}(x-\theta)^{2}$
$=\beta(p+1/2)x^{2}+[\beta(p+1/2)^{2}-\alpha](x-\theta)^{2}+f(x)$
$=Q(x)+f(x)$
$\leq c_{0}+f(x),$
where $c_{0}>0$ because the quadratic $Q(x)$ is strictly positive. We can rewrite the inequality as
$\mathbbm{E}[g(Y)\mid X=x]$ $\leq mf(x)+n$
$m$ $=1$
$n$ $=c_{0},$
where $m,n>0$. It remains to show that $jm<1$. Since $10<\gamma<$ 1 by assumption, the proof is now complete. ∎
Proposition 4.
If $X\sim\mathrm{GIG}(p,a,b)$ and $p>1$, then $X=_{d}Y+E$ for independent random variables $Y$ and $E$ with
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $={K_{p-2}(\sqrt{ab})}/{K_{p}(\sqrt{ab})}\in(0,1)$
and $E\sim\mathrm{Exp}(a/2).$
Proof.
First, we prove that the PDF of the generalized inverse Gaussian distribution for $p>1$ can be written as a continuous mixture of truncated exponential random variables. After that, it will be straightforward to prove the main result. More precisely, we show that
$f_{X}(x)$ $=\int_{0}^{\infty}f_{X\mid Y}(x\mid y)f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{\infty}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\mathbbm{1}(x\geq y)\,f_{Y}(y)\,\mathrm{d}y,$
where
$f_{Y}(y)$ $=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b)$
$w$ $=\frac{K_{p-2}(\sqrt{ab})}{K_{p}(\sqrt{ab})}.$
Let $a>0,b>0,p>1$, and define:
[…]
The ratio
[…]
is monotone non-decreasing in $x>0$ for $p>1$, so we can apply Theorem 2.4 in [20] to write $g(x)$ as a mixture of truncated $\mathrm{Exp}(a/2)$ random variables:
$g(x)=\mathrm{GIG}(x\mid p,a,b)=\int_{0}^{\infty}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\mathbbm{1}(x\geq y)\,f_{Y}(y)\,\mathrm{d}y.$
By Theorem 2.4 in [20], the mixing density $f_{Y}(y)$ is
$f_{Y}(y)$ $=\left(\frac{g(y)}{f(y)}\right)^{\prime}\,\int_{y}^{\infty}f(s)\,\mathrm{d}s$
[…]
$=w\,\mathrm{GIG}(y\mid p-2,a,b)+(1-w)\,\mathrm{GIG}(y\mid p-1,a,b).$
Now, we prove the main result. Let $\tilde{X}=Y+E$. It remains to show that $\tilde{X}=_{d}X$. By the convolution formula,
$f_{\tilde{X}}(x)$ $=\int_{0}^{\infty}f_{E}(x-y)f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{\infty}\mathrm{Exp}(x-y\mid a/2)\,f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{\infty}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\mathbbm{1}(x\geq y)\,f_{Y}(y)\,\mathrm{d}y$
$=\mathrm{GIG}(x\mid p,a,b),$
as we wanted to show. ∎
Proposition 5.
Let $G_{p}(x)$ be the CDF of the $\mathrm{GIG}(p,a,b)$. For half-integer $p$ such that $|p|>1/2$, the following recurrence holds:
$G_{p}(x)=wG_{p-2}(x)+(1-w)G_{p-1}(x)-\exp\left\{\frac{-ax}{2}\right\}I_{p}(x),$
where
$I_{p}(x)=\frac{w(ab)^{(p-2)/2}}{2^{p-1}K_{p-2}(\sqrt{ab})}\Gamma\left(2-p,\frac{b}{2x}\right)+\frac{(1-w)(ab)^{(p-1)/2}}{2^{p}K_{p-1}(\sqrt{ab})}\Gamma\left(1-p,\frac{b}{2x}\right)$
and $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1].
Proof.
Let $G_{p}$ be the CDF of $\mathrm{GIG}(p,a,b)$. Then,
$G_{p}(x)$ $=\int_{0}^{\infty}\left(\int_{y}^{\max(x,y)}\frac{a}{2}\exp\left\{-\frac{a}{2}(x-y)\right\}\,\mathrm{d}x\right)\,f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{\infty}\left[1-\exp\left\{\frac{a(y-\max(x,y))}{2}\right\}\right]\,f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{x}\left[1-\exp\left\{\frac{a(y-x)}{2}\right\}\right]\,f_{Y}(y)\,\mathrm{d}y$
$=\int_{0}^{x}f_{Y}(y)\,\mathrm{d}y-\exp\left(\frac{-ax}{2}\right)\int_{0}^{x}\exp\left(\frac{ay}{2}\right)f_{Y}(y)\,\mathrm{d}y.$
On the one hand,
$\int_{0}^{x}f_{Y}(y)\,\mathrm{d}y=wG_{p-2}(x)+(1-w)G_{p-1}(x),$
where $G_{p-2}$ and $G_{p-1}$ are the CDFs of $\mathrm{GIG}(p-2,a,b)$ and $\mathrm{GIG}(p-1,a,b)$, respectively. On the other hand,
$\int_{0}^{x}\exp\left(\frac{ay}{2}\right)f_{Y}(y)\,\mathrm{d}y=I_{p}(x)$
$I_{p}(x)=\frac{w(ab)^{(p-2)/2}}{2^{p-1}K_{p-2}(\sqrt{ab})}\Gamma\left(2-p,\frac{b}{2x}\right)+\frac{(1-w)(ab)^{(p-1)/2}}{2^{p}K_{p-1}(\sqrt{ab})}\Gamma\left(1-p,\frac{b}{2x}\right),$
where $\Gamma(\cdot,\cdot)$ is the incomplete gamma function [1]. ∎